# Circuit Insights: Towards Interpretability Beyond Activations

### v1

@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption in recent years, including in sensitive domains such as medical analysis (Singhal et al., 2023). Despite their remarkable capabilities, understanding the internal mechanisms of these models is crucial for safe and reliable deployment, yet our knowledge in this area remains limited (Olah et al., 2018; Lapuschkin et al., 2019; Sharkey et al., 2025). Several methods have emerged in the fields of mechanistic interpretability and explainable AI to understand how models encode and utilize mechanisms that influence outputs (Olah et al., 2020; Achtibat et al., 2024; Dreyer et al., 2025). Much of the existing work focuses on circuit discovery, identifying subgraphs responsible for specific tasks (Conmy et al., 2023). However, these studies are mostly limited to toy tasks, and understanding the roles of individual neurons and attention heads still requires extensive manual analysis (Elhage et al., 2022; Wang et al., 2022; Bricken et al., 2023).
@@
Automated interpretability methods have been proposed to address these limitations. Initial work, such as Bills et al. (2023), leveraged larger LLMs to analyze activation patterns of MLP neurons and generate natural language descriptions. Although promising, these approaches face the fundamental challenge of polysemanticity of MLP neurons, making them inherently difficult to interpret. This bottleneck prompted the development of Sparse Autoencoders (SAEs), which decompose activations into more monosemantic features (Bricken et al., 2023), advancing scalable interpretability pipelines (Templeton et al., 2024; Paulo et al., 2025a). More recently, transcoders were introduced in Dunefsky et al. (2024) and Ge et al. (2024) as an alternative approach for extracting sparse features. Unlike SAEs, which reconstruct activations, transcoders sparsely approximate entire MLP layers, while maintaining a clear separation between input-dependent and weight-dependent contributions. This architecture enables efficient circuit discovery and provides direct attributions to other features, attention heads and vocabulary.
@@
Despite these advances in sparse feature space construction, automated interpretability remains heavily dependent on explainer LLMs, which shifts the solution to the black box problem to yet another black box LLM, resulting in notable safety risks which may produce unfaithful or unreliable explanations (Lermen et al., 2025). Its effectiveness is influenced by the prompt, fine-tuning strategy, and the dataset used for generating explanations. Furthermore, sparse features can still be challenging to interpret (Puri et al., 2025), as they may activate on highly specific patterns that are not easily captured by analyzing activations alone, or may be polysemantic.
@@
In this work, we focus on automated interpretability grounded in model weights and circuit structure, making the following contributions:
@@
  • We introduce WeightLens, a framework for interpreting models using only their weights and the weights of their transcoders, reducing dependence on both the underlying dataset and explainer LLMs. Descriptions obtained via WeightLens match or exceed activation-based descriptions for context-independent features.
@@
  • We introduce CircuitLens, a framework for circuit-based analysis of feature activations, extending interpretability to context-dependent features by (i) isolating input patterns triggering feature activations and (ii) identifying which model outputs are influenced by specific features.
@@
Our approach uncovers complex patterns invisible to activation-only methods, addresses the large dataset requirements and explainer LLM dependence of autointerpretability pipelines (Choi et al., 2024; Puri et al., 2025). Additionally, it handles polysemanticity through circuit-based clustering and combining their interpretations into unified feature descriptions.
@@ 2 Related Work
Recently, a series of works has focused on building automated interpretability pipelines for language models (Choi et al., 2024; Paulo et al., 2025a; Puri et al., 2025; Gur-Arieh et al., 2025). Most approaches follow the framework introduced by Bills et al. (2023), which consists of running a large dataset through a model, collecting maximally activating samples for each neuron or SAE feature, applying various sampling strategies, and then passing the samples to a larger LLM to generate natural language descriptions.
Several studies have refined this pipeline by focusing on prompt construction and description evaluation. For instance, Choi et al. (2024) and Puri et al. (2025) examine how factors such as the number of samples and the presentation of token activations affect description quality. Choi et al. (2024) further fine-tune an explainer model specifically to produce descriptions conditioned on a feature’s activations. Beyond input-based evidence, Gur-Arieh et al. (2025) incorporate output-side information, analyzing not only what inputs trigger a feature but also how that feature influences the model’s logits.
@@
Finally, Puri et al. (2025) highlight a persistent challenge: even though sparse features, specifically in SAEs, are generally more monosemantic than MLP neurons, they can be highly specific and activate only on certain patterns. This specificity makes them difficult to interpret; either the explainer LLM fails to identify the correct trigger, or the resulting description is too vague to be useful.333neuronpedia.org/gemma-2-2b/4-gemmascope-transcoder-16k/13598
@@ 3 Methodology
As introduced by Dunefsky et al. (2024), given a transcoder structure, the contribution of transcoder feature $i^{\prime}$ in transcoder layer $l^{\prime}$ to feature $i$ in layer $l>l^{\prime}$ on token $t$ can be expressed as:
@@
3.1 Input-Invariant Analysis
Figure 1: Process of layerwise input-invariant analysis via WeightLens.
The input-invariant connections from Equation (1) provide a useful foundation for interpretation of transcoder features, as demonstrated in the case studies of (Dunefsky et al., 2024). To build on this idea, we make the following complementary assumptions:
Since many features are context-dependent, relying solely on weight-based analysis can produce misleading results. To address this, we introduce a validation step to determine whether a feature is truly token-dependent, i.e., whether it consistently activates on specific tokens regardless of context. Formally, we state the following assumption:
Assumption 2: If a token is strongly supported by input-invariant connections (weights) and semantically aligned with the concept encoded by the feature, then the feature should activate on this token regardless of context.
We generate a feature description – a set of tokens associated with the feature – by processing the model layer-by-layer, starting from layer 0 as presented on Figure 1. For this we are performing the following steps:
  1. 1.
    Extract candidate tokens from vocabulary and previous layers features: Project the feature encoder vector $f_{\text{enc}}$ into the input vocabulary embedding space via the embedding matrix $W_{emb}$ as $W_{emb}\cdot f_{\text{enc}}$, and identify candidate tokens as statistical outliers based on their z-scores (Barnett & Lewis, 1994), retaining only highly distinctive tokens. For each earlier layer $l^{\prime}<l$, compute $W_{\text{dec}}^{(l^{\prime})}\cdot f_{\text{enc}}^{(l)}$, identifying top contributing (outlier) features using the same criterion, and inherit their token descriptions.
  2. 2.
    Validate tokens: Retain only tokens that activate the feature in a forward pass.
  3. 3.
    Analyze output effects: Project the feature decoder vector $f_{\text{dec}}$ into vocabulary logits via $f_{\text{dec}}\cdot W_{U}$, and identify outlier tokens (strongly promoted) via z-score.
Token-based features often respond to multiple forms of the same word. To process the obtained set of tokens and produce a coherent feature description, we apply lemmatization (Bird et al., 2009) to both the promoted tokens and the generated descriptions. This step consolidates different inflected forms into a single canonical form and can be considered a lightweight alternative to LLM-based postprocessing for cleaning and standardizing the descriptions.
3.2 Circuit-Based Analysis
Figure 2: Types of attributions in transcoders.
To account for interference between layers, Ameisen et al. (2025) propose incorporating a Jacobian term into the attribution formulation, which improves the reliability of feature attributions. With this adjustment, Equation (1) can be redefined as
@@ 3.2 Circuit-Based Analysis
Interpretability Beyond Activations A central challenge in interpreting feature activations is that raw activation values do not always reveal what triggered an activation of a given feature. Simply highlighting token activations and prompting language models often yields vague or generic explanations such as “variety of words on variety of topics.”
@@
To address this, we focus on identifying patterns in the data that both lead to a feature’s activation and determine how the feature influences the output.
@@
  • Input-centric focus: Using the attribution formulation in Equation (4), we extract attention head–token pairs that provide strong contribution for a feature’s activation, as presented in Figure 2a. Outlier pairs are selected based on their $z$-score relative to the distribution of contributions, ensuring that only the strongest connections are retained. We then mask the original input sequence, keeping only tokens that either directly activated the feature or contributed significantly through attention. This procedure isolates interpretable token patterns underlying the activation of a feature, as illustrated in Figure 2b, where the feature activates on references to already mentioned entities.
@@
  • Output-centric analysis: Using Equation (5), we evaluate whether the analyzed feature contributed to the prediction of the generated tokens after being activated. This highlights which output tokens were influenced by the feature and thus provides an estimate of its downstream impact, as shown in Figure 2c.
@@
Circuit-Based Clustering A single feature often responds to multiple concepts, which may be entangled and hard to interpret. Semantic clustering of activations in embedding space is insufficient, as it ignores the causal, circuit-level mechanisms.
@@
We propose circuit-based clustering: for each input, we collect contributing elements, such as transcoder features and token/attention head pairs via Equation (2) and Equation (4) respectively, including significant transcoder features $(l^{\prime},i^{\prime})$ and attention head contributions $(l,h,\Delta)$, where $\Delta$ is the relative token position. Sparse activations ensure that each input has only a few contributors.
@@
Sampling Strategy Most prior works (Bills et al., 2023; Choi et al., 2024; Gur-Arieh et al., 2025; Puri et al., 2025) focus on analyzing the most highly activating examples of a feature. This approach is especially sensible for MLP neurons, whose activations are often noisy and unstructured. In contrast, both SAEs and transcoders are explicitly designed to yield monosemantic features. For this reason, we aim to analyze the entire distribution of a feature’s activations, in order to capture the broader concept(s) that drive its behavior.
Automated Interpretability Sampled inputs are first analyzed from input- and/or output-centric perspectives, then grouped into clusters according to the detected circuits. Each cluster is interpreted independently using an explainer LLM (GPT-4o-mini; see Appendix C.2). Rather than providing full inputs with highlighted activations or token–activation pairs, we supply only the detected pattern, marking the single most activating token. Finally, for each feature, the explainer LLM synthesizes a unified description from the individual cluster-level interpretations.
@@ 3.3 Evaluation
All evaluations were performed on input-centric metrics Clarity, Responsiveness, Purity and the output-centric metric Faithfulness from the FADE framework (Puri et al., 2025). Here, Clarity measures if the concept is expressed clearly enough to generate synthetic data, that would activate a feature. Responsiveness demonstrates if the feature’s activation on the given concept are significantly higher than it’s activations on a normally distributed dataset. Purity measures if the feature only strongly activates on the described concept or also on unrelated concepts. Finally, Faithfulness measures the extent to which steering the feature influences the model output in the direction of the described concept.
4 Weight-Based Interpretability Results
We analyze the interpretability of transcoders for GPT-2 Small (Dunefsky et al., 2024), Gemma-2-2b (Lieberum et al., 2024), and Llama-3.2-1B (Paulo et al., 2025b) via WeightLens, and we focus on Gemma-2-2b transcoders for qualitative analysis. We evaluate and compare:
  • WeightLens (Weights-based descriptions): descriptions composed solely of the tokens that activate the feature, postprocessed via lemmatization;
  • WeightLens+Out (Weight-based with logits-based analysis): weight-based descriptions augmented with tokens derived from the feature’s unembedding vectors;
  • WeightLens+Output+LLM (Weight-based with logits and LLM refinement): descriptions further refined using a secondary LLM (gpt-4o-mini-2024-07-18) to generate a concise, descriptive single-line summary based on both activating and promoted tokens (see Appendix C.1).
They are further compared to activation-based methods, specifically descriptions from Neuronpedia (Lin, 2023) and MaxAct* method (Puri et al., 2025).
Layer 0 Layer 7 Layer 12 Layer 21
C R P F C R P F C R P F C R P F
WeightLens 0.56 0.24 0.03 0.00 0.62 0.22 0.02 0.01 0.47 0.13 0.02 0.01 0.71 0.86 0.65 0.02
WeightLens+Out 0.51 0.26 0.03 0.00 0.55 0.24 0.02 0.01 0.39 0.15 0.02 0.01 0.63 0.87 0.63 0.04
WeigthLens+Out+LLM 0.52 0.27 0.04 0.01 0.58 0.20 0.03 0.02 0.41 0.13 0.03 0.00 0.68 0.84 0.63 0.03
Neuronpedia 0.53 0.22 0.07 0.03 0.51 0.17 0.10 0.03 0.28 0.10 0.08 0.00 0.50 0.80 0.73 0.02
MaxAct* 0.50 0.21 0.08 0.04 0.54 0.15 0.08 0.05 0.33 0.11 0.7 0.00 0.53 0.80 0.74 0.03
Table 1: Evaluation of Gemma-2-2b transcoder descriptions (C = Clarity, R = Responsiveness, P = Purity, F = Faithfulness; metrics in range 0–1, higher is better). Best results marked in bold. Methods: WeightLens variants; Circuit-Based; Neuronpedia and MaxAct* (activation-based).

### v2

@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption in recent years, including in sensitive domains such as medical analysis (Singhal et al., 2023). Despite their remarkable capabilities, understanding the internal mechanisms of these models is crucial for safe and reliable deployment, yet our knowledge in this area remains limited (Olah et al., 2018; Lapuschkin et al., 2019; Sharkey et al., 2025). Several methods have emerged in the fields of mechanistic interpretability and explainable AI to understand how models encode and utilize mechanisms that influence outputs (Olah et al., 2020; Achtibat et al., 2024; Dreyer et al., 2025). Much of the existing work focuses on circuit discovery, identifying subgraphs responsible for specific tasks (Conmy et al., 2023). However, these studies are mainly limited to toy tasks, and understanding the roles of individual neurons and attention heads still requires extensive manual analysis (Elhage et al., 2022; Wang et al., 2022; Bricken et al., 2023).
@@
Automated interpretability methods were proposed to address these limitations. Bills et al. (2023) suggested using larger LLMs to analyze activation patterns of MLP neurons and generate natural language descriptions. Although promising, such an approach faces the fundamental challenge of polysemanticity of MLP neurons, making them inherently difficult to interpret. This bottleneck prompted the development of Sparse Autoencoders (SAEs), which decompose activations into monosemantic features (Bricken et al., 2023), advancing interpretability pipelines (Templeton et al., 2024; Paulo et al., 2025a). More recently, transcoders were introduced in Dunefsky et al. (2024) and Ge et al. (2024) as an alternative approach for extracting sparse features. Unlike SAEs, which reconstruct activations, transcoders sparsely approximate entire MLP layers, while maintaining a clear separation between input-dependent and weight-dependent contributions. This architecture enables efficient circuit discovery and provides direct attributions to other features, attention heads and vocabulary.
@@
Despite these advances in sparse feature space construction, automated interpretability remains heavily dependent on explainer LLMs, which shifts the solution to the black box problem to yet another black box LLM, resulting in notable safety risks that may produce unfaithful or unreliable explanations (Lermen et al., 2025). Its effectiveness is influenced by the prompt, fine-tuning strategy, and the dataset used for generating explanations. Furthermore, sparse features can still be challenging to interpret (Puri et al., 2025), as they may activate on highly specific patterns that are not easily captured by analyzing activations alone, or may be polysemantic (Kopf et al., 2025).
@@
In this work, we focus on automated interpretability grounded in model weights and circuit structure. Our contributions consist of the following methods:
@@
  • WeightLens: a framework for interpreting models using their weights and the weights of their transcoders, eliminating dependence on the underlying dataset and explainer LLMs. Descriptions obtained via WeightLens match or exceed baselines for token-based features.
@@
  • CircuitLens: a framework for circuit-based analysis of activations, extending interpretability to context-dependent features by (i) isolating input patterns triggering feature activations and (ii) identifying which output tokens are influenced by specific features.
@@
Circuit-based approach uncovers patterns missed by activation-only methods, simplifying the explainer LLM’s task and improving robustness on smaller datasets by combining weight-based and circuit-based information. Additionally, it handles polysemanticity through circuit-based clustering and combining their interpretations into unified feature descriptions.
@@ 2 Related Work
Recently, a series of works focused on building automated interpretability pipelines for LLMs (Choi et al., 2024; Paulo et al., 2025a; Puri et al., 2025; Gur-Arieh et al., 2025). Those approaches follow the framework introduced by Bills et al. (2023), which consists of passing a large dataset through a model, collecting maximally activating samples for each MLP neuron, and then forwarding those samples with their per-token activations to a larger LLM to generate natural language descriptions.
Several studies refine this pipeline by focusing on prompt construction and description evaluation. For instance, Choi et al. (2024) and Puri et al. (2025) examine how factors such as the number of samples and the presentation of token activations affect description quality. Choi et al. (2024) further fine-tune an explainer model to produce descriptions conditioned on a feature’s activations. Beyond input-based evidence, Gur-Arieh et al. (2025) incorporate output-side information, analyzing not only what inputs trigger a feature but also how that feature influences the model’s logits.
@@
Finally, Puri et al. (2025) highlight a persistent challenge: even though sparse features, specifically in SAEs, are generally more monosemantic than MLP neurons, they can be highly specific and activate only on certain patterns. This specificity makes them difficult to interpret; either the explainer LLM fails to identify the correct trigger, or the resulting description is too vague to be useful.
@@ 3 Methodology
3.1 Attribution in Transcoders
Figure 1: Types of attributions in transcoders.
We leverage circuit-based information for automated interpretability. Due to their architecture, transcoders enable efficient computation of attributions, which we demonstrate in this section.
As introduced by Dunefsky et al. (2024), given a transcoder structure, the attribution of transcoder feature $i^{\prime}$ in transcoder layer $l^{\prime}$ to feature $i$ in layer $l>l^{\prime}$ on token $t$ can be expressed as:
@@ 3.1 Attribution in Transcoders
Ameisen et al. (2025) propose incorporating a Jacobian term into the attribution formulation, which improves the reliability of feature attributions. With this adjustment, Eq. (1) can be redefined as
@@
3.2 WeightLens: Input-Invariant Automated Interpretability
Figure 2: Input-invariant feature analysis with WeightLens. Step 1: Project the feature’s encoder and decoder using the embedding and unembedding matrices. Step 2: Identify top contributing features from earlier layers. Step 3: Validate candidate tokens via a forward pass, keeping only those that activate the feature; use them and the output-based outliers for constructing the description.
The input-invariant connections from Eq. (1) provide a useful foundation for interpretation of transcoder features, as demonstrated by Dunefsky et al. (2024). Qualitatively, we observe that many strong weight-based connections align with the connections revealed by attribution graphs on feature-relevant inputs, as shown in Appendix E.1. These connections help identify which tokens contribute to a feature, either directly through the embeddings matrix or indirectly by activating upstream features. At the same time, it can be difficult to distinguish meaningful connections that persist when an input is provided to the model, from those that are simply noise. To build on this idea, we make the following assumption:
Since many features are context-dependent, weight-based analysis alone can be misleading. Connections identified under Assumption 1 do not always hold across contexts (see Appendix E.3). Since our goal is to identify tokens that reliably characterize a feature, we introduce a validation step to check whether the feature activates on certain tokens in isolation or depends on context, in which case no stable token-level description can be assigned. Formally, we state the assumption:
Assumption 2: If a token is strongly supported by input-invariant connections to the feature (weights) and reflects the feature’s inherent behavior, then the feature should activate on this token regardless of context or in the absence of it.
Based on these assumptions, we generate token-based feature descriptions, where each feature is characterized by tokens that activate it and by tokens it promotes at the output. The procedure is applied layer by layer, progressing from earlier to later layers. The following steps, illustrated in Figure 2, describe the full process.
  1. 1.
    Embedding and unembedding projections: Project the feature encoder vector $f_{\text{enc}}$ into the input embedding space using $W_{emb}\cdot f_{\text{enc}}$, and, following Gur-Arieh et al. (2025), the feature decoder vector $f_{\text{dec}}$ into vocabulary logits via $f_{\text{dec}}\cdot W_{U}$. Detect outliers using $z$-scores (Barnett & Lewis, 1994): embedding-space outliers are candidate tokens for the input-based part of the feature description, and logit-space outliers form the output-based part (outlier threshold 4 for GPT-2 (Radford et al., 2019); 4.5 for Gemma-2-2B (Riviere et al., 2024) and Llama-3.2-1B (Dubey et al., 2024)) .
  2. 2.
    Analyze weight-based feature connections: For each earlier layer $l^{\prime}<l$, compute $W_{\text{dec}}^{(l^{\prime})}\cdot f_{\text{enc}}^{(l)}$ to identify top contributing features (outliers with $z$-score threshold 3). Inherit their token descriptions as candidate tokens for the input-based part of the feature description.
  3. 3.
    Validate candidate tokens and generate description: Keep only tokens that activate the feature in a forward pass. Combine these validated input-based tokens and (optionally) the output-based tokens to form the final feature description. Note that output-based tokens might be noisy and we cannot validate them, therefore this part is optional and its necessity is further evaluated in the Section 4.
Token-based features often respond to multiple forms of the same word. To process the obtained set of tokens and produce a coherent feature description, we apply lemmatization (Bird et al., 2009) to the generated descriptions. This step consolidates different inflected forms into a single canonical form and can be considered a lightweight alternative to LLM-based postprocessing for cleaning and standardizing the descriptions.
3.3 CircuitLens: Automated Interpretability Based on Circuits
In transformer models, much of the behavior depends on context and cannot be interpreted from weights alone. Following Bills et al. (2023), most automated interpretability methods select samples that strongly activate a feature across a large dataset and pass them, along with per-token activations, to an explainer LLM to generate a description. However, activations do not always reveal which input tokens caused the feature to fire, and noisy or polysemantic inputs can further reduce description quality. CircuitLens addresses these limitations by providing a circuit-aware approach to feature interpretability. Below, we describe the whole process of automated interpretability with CircuitLens.
Activation Caching and Sampling Most prior works (Bills et al., 2023; Choi et al., 2024; Gur-Arieh et al., 2025; Puri et al., 2025) focus on analyzing the most highly activating examples of a feature. This approach is especially sensible for MLP neurons, whose activations are often noisy and unstructured. In contrast, both SAEs and transcoders are explicitly designed to yield monosemantic features. For this reason, we aim to analyze the entire distribution of a feature’s activations, in order to capture the broader concept(s) that drive its behavior. For this, we cache maximum sample activations over the whole dataset, which is feasible due to high sparsity of features.
Circuit-Based Patterns Detection
@@ 3.3 CircuitLens: Automated Interpretability Based on Circuits
A central challenge in interpreting feature activations is that raw activation values do not always reveal what triggered an activation of a given feature. Simply highlighting token activations and prompting language models often yields vague or generic explanations such as “variety of words on variety of topics.”333neuronpedia.org/gemma-2-2b/4-gemmascope-transcoder-16k/13598
@@
To address this, we focus on identifying patterns in the data that both lead to a feature’s activation and determine what parts of the output, produced by the model, was influenced by the feature.
@@
  • Input-centric focus: Using the attribution formulation in Eq. (4), we extract (attention head, token) pairs that provide strong contribution for a feature’s activation. Outlier pairs are selected based on their $z$-score relative to the distribution of contributions, ensuring that only the strongest connections are retained. We then mask the original input sequence, keeping only tokens that either directly activated the feature or contributed significantly through attention. This procedure isolates interpretable token patterns underlying the activation of a feature, as illustrated in Figure 1b, where the feature activates on references to already mentioned entities.
@@
  • Output-centric analysis: Using Eq. (5), we evaluate whether the analyzed feature contributed to the prediction of the generated tokens after being activated. This highlights which output tokens were influenced by the feature and thus provides an estimate of its downstream impact, as shown in Figure 1c.
By masking only the input tokens that activated the feature and highlighting only the output tokens it influenced, we pre-identify the relevant patterns ourselves. This removes the burden from the explainer LLM of searching through the full context and instead lets it focus on describing the common concept behind these selected inputs and outputs, hence simplifying its task.
@@
Circuit-Based Clustering In order to simplify the explanation task further, we want the explainer LLM to receive clear, monosemantic inputs. However, a single feature can respond to multiple entangled concepts, which remain hard to interpret even in SAEs (Kopf et al., 2025). Semantic clustering of activations in embedding space is not always sufficient, since it ignores the underlying causal, circuit-level mechanisms.
@@
We propose circuit-based clustering: for each input, we collect contributing elements, such as transcoder features and token/attention head pairs via Eq. (2) and Eq. (4) respectively. Significant features $(l^{\prime},i^{\prime})$ and attention head contributions $(l,h,\Delta)$, where $\Delta$ is the relative token position, are combined into a single vector. Sparse activations ensure that each input has only a few contributors.
@@
Generating Descriptions Sampled inputs are first analyzed from input- and/or output-centric perspectives, then grouped into clusters according to the detected circuits. Each cluster is interpreted independently using an explainer LLM (see Appendix F.2 for prompts). Rather than providing full inputs with highlighted activations or token–activation pairs, we supply only the detected pattern, marking the single most activating token. Finally, for each feature, the explainer LLM synthesizes a unified description from the individual cluster-level interpretations.
@@ 3.4 Evaluation
All evaluations use the FADE framework (Puri et al., 2025), which relies on an LLM-as-a-judge to assess alignment between a feature and its natural-language description across four metrics: Clarity, Responsiveness, Purity, and Faithfulness. This evaluation setup follows prior work showing strong correlations between such automated judgments and human assessments of description quality.
Clarity measures whether the description expresses the concept clearly enough for the explainer LLM to generate synthetic inputs that activate the feature. Responsiveness checks whether the feature activates more strongly on these concept-driven inputs than on a randomly sampled dataset. Purity measures whether the feature activates primarily on inputs that match the described concept, or also on unrelated inputs. Faithfulness evaluates how strongly the feature’s activation influences the model’s output in the direction implied by the description. A review of automated interpretability metrics, along with details of FADE’s computation procedure, is provided in Appendix A.2.
4 WeightLens: Evaluations and Findings
We analyze the interpretability of transcoders for GPT-2 Small (Dunefsky et al., 2024), Gemma-2-2B (Lieberum et al., 2024), and Llama-3.2-1B (Paulo et al., 2025b), focusing on Gemma-2-2B for qualitative analysis. We evaluate and compare the following setups:
  • WeightLens: descriptions composed solely of validated tokens that activate the feature.
  • WeightLens+Out: WeightLens descriptions augmented with tokens promoted by the feature, derived from its unembedding projection.
  • WeightLens+Out+LLM: descriptions refined using an explainer LLM (gpt-4o-mini-2024-07-18) instead of lemmatization, generating a concise single-line summary based on both activating and promoted tokens (see Appendix F.1 for details).
With these setups, we aim to demonstrate the effect of including embedding-based contributions, output-based signals, and LLM refinement on feature description quality. They are further compared to activation-based methods, specifically descriptions available on Neuronpedia (Lin, 2023), as well as generated via MaxAct* method presented by Puri et al. (2025), details are in the Appendix D.
Figure 3: Evaluation of Gemma-2-2B transcoder descriptions (C = Clarity, R = Responsiveness, P = Purity, F = Faithfulness). Methods: WeightLens variants. Baselines: Neuronpedia and MaxAct*.